# Privacy-Aware Semantic Cache for Large Language Models

### v1

@@ 3.2. MeanCache’s System Design
MeanCache is a user-centric semantic caching system, optimized to function on the user side. It enhances the performance of the semantic caching system by leveraging the collective intelligence of multiple users. The system design of MeanCache is briefly described as follows: (1) MeanCache employs a small embedding model, which has a lower computational overhead compared to large Language Learning Model (LLM) based embedding models, such as Llama 2. (2) MeanCache uses federated learning for training to fine-tune the embedding model. This approach generates high-quality embeddings and improves the accuracy of embedding matching for retrieving similar queries. Notably, federated learning does not necessitate the storage of user queries in a central location, as it can train a high-quality model using the users’ distributed data. (3) To reduce storage and memory overhead, as well as to expedite the search time for finding similar queries in the cache, MeanCache compresses the embeddings using Principal Component Analysis (PCA). (4) MeanCache autonomously determines the optimal threshold for cosine similarity. It does this by varying the threshold and selecting the one that optimizes the F-score of the cache. Next we describe the system design of MeanCache in detail.
3.2.1. Privacy-Preserving Embedding Model FL Training
@@
FL training is a decentralized approach to train machine learning models. It allows the users/clients to train the embedding model locally and share the updated weights with the server. The server aggregates the updated weights from multiple users, evaluates the global model, and redistributes the updated model to the users. This approach ensures that the user’s privacy is maintained, and the collective intelligence of multiple users is leveraged to improve the performance of the caching system. In the first step, the server sends the initial weights of the embedding model ($W_{global}^{t+1}$) to a subset of users. The subset of users are randomly selected or selected based on their battery level, network bandwidth, or performance history. Additionally, the server also sends the hyperparameters (e.g., learning rate, batch size, epochs) necessary for FL training of the embedding model.
Client Training.
@@
Upon receiving the embedding model ($W_{global}^{t+1}$) from the server, each client replaces its local embedding model weights with the newly received weights. Subsequently, each client proceeds to train the embedding model locally on its unique local dataset. Local training employs a multitask learning approach, integrating two distinct loss functions: contrastive loss  (Reimers and Gurevych, 2019) and multiple-negatives ranking loss  (Henderson et al., 2017; Reimers and Gurevych, 2019). These loss functions update the weights of the embedding model during the training process. The contrastive loss function operates by distancing unique (non-duplicate) queries from each other within the embedding space, thereby facilitating the differentiation between duplicate and non-duplicate queries. Unlike contrastive loss, the multiple-negatives ranking loss function minimizes the distance between positive pairs (duplicate queries) amidst a large set of potential candidates i.e., multiple-negatives ranking loss does not concentrate on distancing unique queries and its primary objective is to draw positive pairs (similar queries) closer within the embedding space.
@@
This multitasks learning approach enables MeanCache to adjust to diverse query patterns exhibited by users. For instance, some users may generate more repetitive queries compared to others, while certain users may not produce any repetitive queries at all. Interestingly, MeanCache’s multitask learning objective can benefit from learning even from a user with no repetitive queries. This is because MeanCache’s global embedding model ($W_{global}^{t+1}$) will learn to widen the distance between unique queries, thereby effectively learning true misses of the non-duplicate queries and consequently minimizing the false hits during the search process.
Finding the Optimal Threshold for Cosine Similarity.
@@
In addition to privacy preserving training of the embedding model, identifying the optimal threshold ($\tau$) for cosine similarity is equally important. The range of $\tau$ is between 0 and 1. This threshold ($\tau$) dictates the level of similarity above which a cached query is considered relevant to the current user query. Setting the threshold too low could result in numerous false hits, leading to the retrieval of irrelevant queries from the cache. Conversely, a threshold set too high might cause many false miss, where relevant queries are not retrieved from the cache. Hence, finding the optimal threshold ($\tau$) for cosine similarity is vital.
@@
After the client’s local training, MeanCache automatically determines this optimal threshold ($\tau$) from the clients’ data. MeanCache’s varies the threshold $\tau$ to find the optimal threshold that optimizes the F-score of the cache (Section 4.2). The client/user will use the optimal threshold to find similar queries in the cache. By finding the optimal threshold, MeanCache effectively balances between true hits and true misses, therefore yielding improved accuracy in semantic similarity matching to return the response from the cache on duplicate queries.
Aggregation.
@@
After client local training and finding the optimal threshold, each client sends updated weights of the global model ($W_{global}^{t+1}$) and optimal threshold ($\tau$) to the server. The server aggregates the updated weights from multiple users to form a new embedding model ($W_{global}^{t+1}$) using FedAvg (McMahan et al., 2017). The server also computes the mean of the received optimal thresholds from the clients for a global optimal threshold ($\tau_{global}$). The benefit of finding $\tau_{global}$ is that when a new user joins the system, the user will not have queries to find its own optimal threshold. In such cases, the system can use $\tau_{global}$ as a starting point for semantic matching.
@@
After finding the global optimal threshold and the global embedding model, the server then redistributes the updated embedding model to the users for the next round of FL training. This process is repeated until the semantic matching accuracy of the MeanCache meets a certain performance threshold (i.e., false hits and false misses are minimized). This approach ensures that the user’s privacy is maintained (i.e., clients queries are never stored at the central server), and the collective intelligence of multiple users is leveraged to improve the performance of the MeanCache. After the completion of the FL training, each client will have access to a fin-tuned embedding model ($W_{global}^{t+1}$) to generate high-quality embeddings which can capture the complex semantics of a user query. For instance, if previously the embedding model was unable to differentiate between ”I love cats” and I do not love cats”, the updated embedding model will be able to differentiate between such queries as it may learn such different semantics from the other users.
3.2.2. Embeddings Compression uisg PCA
@@
MeanCache’s search time performance directly depends on the dimensionality of the embeddings generated by the embedding model. High-dimensional embeddings not only require more memory and storage but also increase the search time for finding similar queries in the cache. This is because the search time is proportional to the dimensions of the embedding vector. Computing the cosine similarity between two high-dimensional embeddings (i.e., new query embeddings and with each embedding in the cache) is computationally expensive and time-consuming. Thus, to improve the search time for finding similar queries in the cache, it is essential to reduce the dimensionality of the embeddings with minimal impact on MeanCache’s performance.
@@
Principal Component Analysis (PCA) is a dimensionality reduction technique that is widely used to compress high-dimensional data into a lower-dimensional space  (Pearson, 1901; Hotelling, 1933; Jolliffe and Cadima, 2016), while still maintaining the most important information. First, MeanCache generates embeddings for all the users’ queries using the embedding model. Next, MeanCache applies PCA to learn the principal components of all the queries embeddings generated in the previous step. MeanCache integrates the learned principal components as an additional layer in the embedding model. This new layer will project the original embeddings onto the lower dimensional space, producing compressed embeddings.
@@
Now, when a non-duplicate query is received, MeanCache will use the updated embedding model (with PCA layer) to generate the compressed embeddings for the new query and store the query, response and the compressed embedding in the cache. Storing the compressed embeddings in the cache will significantly reduce the storage and memory overhead of the embeddings. Next, when a duplicate query is received, MeanCache will use the same embedding model with PCA components to generate the compressed embeddings for this duplicate query and find the similar queries in the cache. Since the embeddings are compressed, the search time for finding similar queries in the cache will be significantly reduced.
@@ 4. Evaluation
We conduct evaluations of MeanCache against the baseline GPTCache (Bang, 2023) to demonstrate that MeanCache achieves optimal performance while preserving user-privacy (i.e., without storing the user queries at the server). For a fair comparison between MeanCache and GPTCache, we employ the optimal configuration as described in the GPTCache study (Bang, 2023). This configuration utilizes Albert (Lan et al., 2019) and sets the cosine similarity threshold at 0.7 to determine the cache hit or miss.
@@
We evaluate MeanCache using the GPTCache dataset. The dataset is partitioned into training, testing, and validation subsets. The training and validation datasets are randomly distributed among the clients, with each client receiving non-overlapping data points. During local training, each client utilizes its training dataset to update its local embedding model and employs the validation dataset to determine the optimal threshold for cosine similarity (Section 4.2). The testing dataset, located at the central server, facilitates a fair comparison between MeanCache and GPTCache (Bang, 2023). It is important to note that usually in real-world deployments, the test dataset may not be available at the central server, and the metrics reported by clients are used to evaluate the performance of the global model.
@@ 4.0.3. Evaluation Metrics
[…]
@@
4.0.4. Research Questions
  • Besides preserving user query privacy, how does MeanCache perform in comparison to GPTCache in terms of performance metrics?
@@ 4.1. Privacy Preserving Embeddings Model Training
Figures 1 and 2 depict the performance of MeanCache as FL training progresses. The X-axis represents the training round, while the Y-axis shows the performance metrics such as F-score, precision, recall, and accuracy of the global model ($W_{global}^{t+1}$). As illustrated in Figure 1, the F-score for MPNet increases from 0.82 to 0.88, and for Albert, it rises from 0.83 to 0.86, as shown in Figure 2. Similarly, precision for MPNet significantly increases from 0.74 to 0.85, as depicted in Figure 1, and for Albert, it increases from 0.74 to 0.81, as demonstrated in Figure 2. Given that MPNet is a more robust transformer architecture compared to Albert, it is also observed during our training that MPNet outperforms Albert, exhibiting superior learning in FL settings.
@@ 4.2. Cosine Similarity Threshold Impact on Semantic Matching
Figures 3 and 4 show how the cosine similarity threshold ($\tau$) affects MeanCache’s performance. The X-axis represents the threshold $\tau$ values, and the Y-axis denotes the performance metrics. For instance, at a 0.3 threshold, MeanCache’s semantic matching accuracy using MPNet is 57%, with a precision of 54% as shown in Figure 3. Similarly, with Albert at the same threshold, the accuracy is 55%, and precision is 53% (Figure 4). Precision typically improves with an increase in threshold. However, beyond a certain point, increasing the threshold $\tau$ leads to a decline in F score, accuracy, and recall, i.e., higher value of $\tau$ increases false cache misses (false negatives). Therefore, setting an optimal threshold is imperative.
@@
Figure 7. Confusion matrices for MeanCache and GPTCache for comprehensive insight. Total queries are 1000, comprising 700 unique and 300 duplicate queries. Among the 700 unique queries, MeanCache produces only 89 false hits, while GPTCache generates a significantly larger number of false hits, 233 in total. False hits are undesirable as they require the user to resend the query to the LLM service to obtain the correct response.
@@ 4.3. MeanCache Comparison with Baseline (GPTCache)
However, Figure 6 shows that GPTCache produces significant false hits on non-duplicate queries (queries 0 to 69) compared to MeanCache, detracting from user experience. Each false hit means the user receives an incorrect query response from the cache and must resend the same query to LLM service, resulting in a poor user experience. To prioritize precision, we adjust the beta value in the F score to 0.5, valuing precision twice as highly as recall to ensure user satisfaction by avoiding false positives. This decision is driven by the need to minimize user inconvenience caused by incorrect cache hits. A false miss (low recall) does not require user intervention as the false miss query will be automatically routed to the LLM. Thus, precision in semantic caching is more important than recall.

### v2

@@ 4. MeanCache’s Design
The key contribution of MeanCache comes from harnessing the collective intelligence of multiple users to training a semantic similarity model and its user-centric design that addresses both privacy and scalability issues. To achieve these, MeanCache takes the following design decisions. It employs a small embedding model with lower computational overhead than large Language Learning Model (LLM) based embedding models, such as Llama 2. It uses federated learning (FL) for collaborative training to fine-tune the embedding model without ever storing user data on a central server. This approach generates high-quality embeddings and improves the accuracy of embedding matching for retrieving similar queries. To reduce storage and memory overhead and expedite the search time for finding similar queries in the cache, MeanCache compresses the embeddings using Principal Component Analysis (PCA). MeanCache leverages user feedback (i.e., forcing LLM request instead of the cache query) to adapt the threshold for cosine similarity. It then shares the updated threshold via FL to determine the optimal threshold across multiple users. It does this by varying the threshold and selecting the one that optimizes the F-score of the cache on clients.
Figure 3. Privacy-Preserving Embedding Model FL Training in MeanCache.
4.1. FL Based Embedding Model Training
GPTCache(Bang, 2023) suggests using Llama 2 to generate superior embeddings, thereby enhancing semantic matching accuracy. However, this approach has several limitations. Large language models like Llama 2 are not only sizable, being gigabytes in size, but they also require substantial computational resources to generate embeddings. Deploying such models, especially at the end-user level, is impractical due to their size and significant computational overhead for semantic matching. MeanCache employs a compact embedding model, which has a lower computational overhead compared to large embedding models. The smaller model may not provide the same level of accuracy as an LLM. However, a smaller model trained on customized tasks can match the performance of an LLM (Ouyang et al., 2022; Penedo et al., 2024; Du and Kaelbling, 2024). One challenge in using smaller embedding models is that each user may not have sufficient data to train and customize the embedding model.
MeanCache utilizes federated learning to exploit the vast amount of distributed data available on users’ devices to train and personalize the smaller embedding model. Federated Learning allows the users to train the embedding model locally and learn the optimal threshold for cosine similarity. The updated weights and local threshold are shared with the server. The server aggregates the updated weights and cosine similarity threshold from multiple users to update the global model, which is redistributed back to the users. This approach ensures that the user’s privacy is maintained, and the collective intelligence of multiple users is leveraged to improve the performance of the caching system. Figure 3 shows the overview of privacy-preserving training of the embedding model with FL.
@@ 4.1. FL Based Embedding Model Training
In the first step, the server sends the initial weights of the embedding model ($W_{global}^{t+1}$) and global threshold ($\tau$) to a subset of users as shown in step 1 in Figure 3. The subset of users is usually randomly selected or selected based on their battery level, network bandwidth, or performance history. Additionally, the server also sends the hyperparameters (e.g., learning rate, batch size, epochs) necessary for FL training of the embedding model.
4.1.1. Client Training.
Upon receiving the embedding model ($W_{global}^{t+1}$) from the server, each client replaces its local embedding model weights with the newly received weights. Subsequently, each client trains the embedding model locally on its unique local dataset (step 2 in Figure 3).
@@
To generate high-quality embeddings from unique and similar queries, MeanCache’s clients’ training employs a multitask learning approach, integrating two distinct loss functions: contrastive loss  (Reimers and Gurevych, 2019) and multiple-negatives ranking loss  (Henderson et al., 2017; Reimers and Gurevych, 2019). These loss functions update the weights of the embedding model during the training process. The contrastive loss function operates by distancing unique (non-duplicate) queries from each other within the embedding space, thereby facilitating the differentiation between duplicate and non-duplicate queries. Unlike contrastive loss, the multiple-negatives ranking loss function minimizes the distance between positive pairs (duplicate queries) amidst a large set of potential candidates i.e., multiple-negatives ranking loss does not concentrate on distancing unique queries and its primary objective is to draw positive pairs (similar queries) closer within the embedding space.
@@
This multitasks learning approach enables MeanCache to adjust to diverse query patterns exhibited by users. For instance, some users may generate more repetitive queries compared to others, while certain users may not produce any repetitive queries at all. Interestingly, MeanCache’s multitask learning objective can benefit from learning even from a user with no repetitive queries. This is because MeanCache’s global embedding model ($W_{global}^{t+1}$) will learn to widen the distance between unique queries, thereby effectively learning the true misses of the non-duplicate queries and minimizing the false hits during the search process. True miss happens when a similar query is not present in the cache. A false hit is when a query is found and returned from the cache, which is not actually similar.
4.1.2. Finding the Optimal Threshold for Cosine Similarity.
After generating query embeddings using an embedding model, a similarity metric such as cosine similarity is used to determine if the new query embeddings match the cached embeddings of past queries. This process involves setting a threshold for cosine similarity, which is a delicate balance.
@@
In addition to privacy-preserving training of the embedding model, MeanCache also learns the optimal threshold ($\tau$) for cosine similarity. The range of $\tau$ is between 0 and 1. This threshold ($\tau$) dictates the level of similarity above which a cached query is considered relevant to the current user query. Setting the threshold too low could result in numerous false hits, leading to retrieving irrelevant queries from the cache. Conversely, a threshold that is set too high might cause many false misses, where relevant queries are not retrieved from the cache. Hence, finding the optimal threshold ($\tau$) for cosine similarity is vital.
@@
During the client’s local training, MeanCache determines this optimal threshold ($\tau$) from the client’s response to the cache query response. Even after finding a cached response, a user requests a response from the LLM, MeanCache considers it as a false positive and adjusts its threshold. MeanCache varies the threshold $\tau$ to find the optimal threshold that optimizes the F-score of the cache (Section 5.3). The user will use the optimal threshold to find similar queries in the cache. By finding the optimal threshold, MeanCache effectively balances between true hits and true misses, therefore yielding improved accuracy in semantic similarity matching to return the response from the cache on duplicate queries.
4.1.3. Aggregation.
@@
After client local training and finding the optimal threshold, each client sends updated weights of the global model ($W_{global}^{t+1}$) and optimal threshold ($\tau$) to the server (step 3 in Figure 3). The server aggregates the updated weights from multiple users to form a new embedding model ($W_{global}^{t+1}$) using FedAvg (McMahan et al., 2017) as shown in step 4 of Figure 3. The server also computes the mean of the received optimal thresholds from the clients for a global optimal threshold ($\tau_{global}$). The benefit of finding $\tau_{global}$ is that when a new user joins the system, the user will not have queries to find its own optimal threshold. In such cases, the system can use $\tau_{global}$ as a starting point for semantic similarity.
@@
After finding the global optimal threshold and the global embedding model, the server then redistributes the updated embedding model to the users for the next round of FL training. This process is repeated over several to improve the semantic matching accuracy of the MeanCache (i.e., lower false hits and false misses). This approach ensures that the user’s privacy is maintained (i.e., clients queries are never stored at the central server), and the collective intelligence of multiple users is leveraged to improve the performance of the MeanCache. After the completion of the FL training, each client will have access to a fine-tuned embedding model ($W_{global}^{t+1}$) to generate high-quality embeddings which can capture the complex semantics of a user query. For example, MeanCache’s embedding model can discern the semantic similarity between ”Draw a line plot in Python” and ”Plot a linear graph using Python”, and serve such queries locally, as depicted in Figure 2.
Figure 4. Embeddings Compression using PCA
4.1.4. Embeddings Compression using PCA
@@
The substantial size of the embedding vector, for instance, the Llama 2 embeddings dimension being 4096, can lead to considerable overhead during the matching process of a new user query embeddings with cached queries embeddings. This is due to the search time being directly proportional to the dimensions of the embedding vector. Furthermore, high-dimensional embeddings demand more memory and storage. For example, the embeddings generated by Llama 2 for a single query require approximately 32.05 KB of memory storage. Consequently, calculating the cosine similarity between two high-dimensional embeddings, specifically new query embeddings and each embedding in the cache, becomes computationally demanding and time-intensive. Therefore, to improve the search time for identifying similar queries in the cache and to reduce the client’s storage needs, it is essential to diminish the dimensionality of the embeddings while ensuring minimal impact on the tool’s performance.
@@
Principal Component Analysis (PCA) is a dimensionality reduction technique that is widely used to compress high-dimensional data into a lower-dimensional space  (Pearson, 1901; Hotelling, 1933; Jolliffe and Cadima, 2016), while still maintaining the most important information. First, MeanCache generates embeddings for all the users’ queries using the embedding model. Next, MeanCache applies PCA to learn the principal components of all the queries embeddings generated in the previous step, as shown in Figure 4-a. MeanCache integrates the learned principal components as an additional layer in the embedding model. This new layer will project the original embeddings onto the lower dimensional space, producing compressed embeddings (Figure 4-b).
@@
When a non-duplicate query is received, MeanCache uses the updated embedding model (with PCA layer) to generate the compressed embeddings (Figure 4-b) for the new query and store the query, response, and the compressed embedding in the cache. Storing the compressed embeddings in the cache will significantly reduce the storage and memory overhead of the embeddings. Next, when a duplicate query is received, MeanCache uses the same embedding model with PCA components to generate the compressed embeddings for this duplicate query and find similar queries in the cache. Since the embeddings are compressed, the search time for finding similar queries in the cache will be significantly reduced.
4.2. Cache Population and MeanCache Implementation
Once the embedding model is trained within MeanCache on each user, it is deployed as depicted in Figure 2. Initially, when a new user starts using MeanCache, the local cache is vacant. During these interactions, if a user query’s response is not found in the cache, the request is forwarded to the LLM web service to retrieve the response, which is then inserted in the cache. If MeanCache finds a semantically similar query for any of the following queries, it responds to the user’s similar queries by retrieving the response from the local cache.
MeanCache is a python-based application that is built on the Flower FL framework (Beutel et al., 2020). A user can submit LLM queries via this application to take advantage of the local cache. The central server, which orchestrates the FL training, may reside on the LLM web service. We employ the Sbert (Reimers and Gurevych, 2019) library to train MPNet (Song et al., 2020) and Albert (Lan et al., 2019) on each client and to generate query embeddings. To efficiently execute a cosine similarity search between a query embedding and cached embeddings, we utilize Sbert’s semantic search, which can handle up to 1 million entries in the cache. MeanCache cache storage is persistent and built using DiskCache (Dis, [n.d.]) library.
@@ 5. Evaluation
Our evaluation answer the following research questions.
  • Besides preserving user query privacy, how does MeanCache perform in comparison to baseline in terms of performance metrics?
5.1. Evaluation Settings
We conduct evaluations of MeanCache against the baseline (Bang, 2023) to demonstrate that MeanCache achieves optimal performance while preserving user-privacy (i.e., without storing the user queries at the server). For a fair comparison between MeanCache and baseline, we employ the optimal configuration as described in the GPTCache study (Bang, 2023). This configuration utilizes Albert (Lan et al., 2019) and sets the cosine similarity threshold at 0.7 to determine the cache hit or miss.
@@ 5.1. Evaluation Settings
We evaluate MeanCache using the GPTCache dataset. The dataset is partitioned into training, testing, and validation subsets. The training and validation datasets are randomly distributed among the clients, with each client receiving non-overlapping data points. During local training, each client utilizes its training dataset to update its local embedding model and employs the validation dataset to determine the optimal threshold for cosine similarity (Section 5.3). The testing dataset, located at the central server, facilitates a fair comparison between MeanCache and GPTCache (Bang, 2023).
@@ 5.1.3. Evaluation Metrics
Fβ Score. A weighted harmonic mean of precision and recall, balancing the two based on the value of $\beta$. $\beta>1$ gives more weight to recall, while $\beta<1$ emphasizes precision.
@@ 5.2. Privacy Preserving Embeddings Model Training
Figures 5 and 6 depict the performance of MeanCache as FL training progresses. The X-axis represents the training round, while the Y-axis shows the performance metrics such as F-score, precision, recall, and accuracy of the global model ($W_{global}^{t+1}$). As illustrated in Figure 5, the F-score for MPNet increases from 0.82 to 0.88, and for Albert, it rises from 0.83 to 0.86, as shown in Figure 6. Similarly, precision for MPNet significantly increases from 0.74 to 0.85, as depicted in Figure 5, and for Albert, it increases from 0.74 to 0.81, as demonstrated in Figure 6. Given that MPNet is a more robust transformer architecture compared to Albert, it is also observed during our training that MPNet outperforms Albert, exhibiting superior learning in FL settings.
@@ 5.3. Cosine Similarity Threshold Impact on Semantic Matching
Figures 7 and 8 show how the cosine similarity threshold ($\tau$) affects MeanCache’s performance. The X-axis represents the threshold $\tau$ values, and the Y-axis denotes the performance metrics. For instance, at a 0.3 threshold, MeanCache’s semantic matching accuracy using MPNet is 57%, with a precision of 54% as shown in Figure 7. Similarly, with Albert at the same threshold, the accuracy is 55%, and precision is 53% (Figure 8). Precision typically improves with an increase in threshold. However, beyond a certain point, increasing the threshold $\tau$ leads to a decline in F score, accuracy, and recall, i.e., higher value of $\tau$ increases false cache misses (false negatives). Therefore, setting an optimal threshold is imperative.
@@
Figure 11. Confusion matrices for MeanCache and GPTCache on 1000 queries comprising 700 unique and 300 duplicate queries. Among the 700 unique queries, MeanCache produces only 89 false hits, while GPTCache generates a significantly larger number of false hits, 233 in total. False hits are undesirable as they require the user to resend the query to the LLM service to obtain the correct response.
@@ 5.4. MeanCache Comparison with Baseline (GPTCache)
However, Figure 10 shows that GPTCache produces significant false hits on non-duplicate queries (queries 0 to 69) compared to MeanCache. Each false hit means the user receives an incorrect query response from the cache and must resend the same query to LLM service, resulting in a poor user experience. To prioritize precision, we adjust the beta value in the F score to 0.5, valuing precision twice as highly as recall to ensure user satisfaction by avoiding false positives. This decision is driven by the need to minimize user inconvenience caused by incorrect cache hits. A false miss (low recall) does not require user intervention as the false miss query will be automatically routed to the LLM. Thus, precision in semantic caching is more important than recall.